OpenStudy (drummergirl7):

Which expressions are equivalent to 3(2x−4)+4x−23(2x−4)+4x−2 ? Select each correct answer. 10x−1410x−14 2(5x−7)2(5x−7) 6x+4x−4−26x+4x−4−2 14x−6

OpenStudy (drummergirl7):

@jabez177

OpenStudy (drummergirl7):

wait lemmie fix this question

OpenStudy (dabdabdab3):

heyallfollowmeoninstagram-taznasser

OpenStudy (drummergirl7):

This is the real question

Directrix (directrix):

Use the Distributive Property 3(2x - 4) + 4x - 2 = 6x -12 + 4x - 2 = ? What do you get when you combine like terms? @DrummerGirl7

OpenStudy (drummergirl7):

Im sorry but i have no clue :(

Directrix (directrix):

6x -12 + 4x - 2 = 6x + 4x - 12 - 2 = 10x - 14 That is one of two correct answers. What is the other correct answer in the list?

OpenStudy (drummergirl7):

14x - 6

Directrix (directrix):

No. 10x - 14 = 2(5x - 7)

OpenStudy (dabdabdab3):

A